# Project Riley: Multimodal Multi-Agent LLM Collaboration with Emotional Reasoning and Voting

### v1

@@ 5.1 Response Comparison: RAG vs No RAG
The first prompt consisted solely of the user’s query: ”Where is the fire happening?”. Responses obtained with and without RAG are presented in LABEL:tab:rag_where.
@@
Analysing LABEL:tab:rag_where, the non-RAG model lacked specific contextual information regarding the ongoing incident, resulting in a generic response containing broad emergency guidance without specifying the fire’s location. Conversely, the RAG-enhanced model delivered a precise answer, clearly identifying the fire’s location as the third floor of a residential apartment at Rua de São Bento 112, Lisbon, Portugal, including coordinates (38.7169° N, 9.1399° W), evacuation status, and detailed safety instructions. This demonstrates the significant advantage of employing RAG for delivering contextually relevant and actionable information in emergency scenarios.
In the second prompt, the user asked: “A fire is happening! What should I do?” with the context “It’s difficult to breathe.”. The responses obtained are detailed in LABEL:tab:rag_breathing.
@@
From LABEL:tab:rag_breathing, both RAG and non-RAG responses effectively addressed the user’s situation by providing clear, practical evacuation and safety instructions suitable for scenarios involving breathing difficulties due to smoke inhalation.
Finally, the third prompt included the user’s query: ”Which number should I call?” with the context ”I’m alone.” and an accompanying image depicting a fire. The responses are summarised in LABEL:tab:rag_call_number.
@@
As observed in LABEL:tab:rag_call_number, the response without RAG provided generic emergency guidance and suggested calling 911 or another local emergency number without explicitly identifying the correct emergency number for Portugal. The RAG-enhanced response correctly identified the emergency number applicable in Portugal (112) but included advice for the user to verify their current country. This suggestion may inadvertently give the impression that the system is uncertain about the incident’s location, despite clearly having access to location-specific data. Therefore, while the RAG implementation successfully provided accurate and relevant information, the prompt formulation could be further refined to avoid potential confusion regarding the system’s knowledge about the user’s geographical context.
@@ 6 Discussion
In the Emotional Appropriateness category, the highest scores occurred in the themes of Job Loss or Unemployment and Difficult Personal Decisions, specifically for the question ”Did the final response seem emotionally appropriate to the context of your question?”, with both themes achieving a mean score of 4.71. This indicates the system’s capability to effectively match emotional tone and content in structured or goal-focused scenarios. Conversely, responses concerning Breakups or Friendship Loss consistently yielded the lowest mean scores in empathy (4.12) and appropriateness (4.41), highlighting challenges in capturing the complexity and variability inherent to interpersonal emotional dynamics.
This trend was further supported by the alignment between user-identified emotions and those suggested by the system, evaluated via the question ”Do you believe the answer(s) of the emotion(s) with the most votes were the most appropriate?”. The strongest alignment appeared in Family or Intergenerational Conflicts (mean score: 4.35), indicating effective resonance with user expectations. In contrast, Breakups or Friendship Loss scored the lowest (mean score: 3.94), suggesting a need for improved modelling of emotional subtleties in relational contexts.
Regarding Clarity and Utility, results were generally positive. Clarity received high ratings across all themes, particularly for Job Loss or Unemployment, with a maximum score of 4.71 for the question ”Was the final response clear and understandable?”. However, perceived utility demonstrated greater variation. The lowest utility scores were associated with Breakups or Friendship Loss (3.88), whereas Difficult Personal Decisions and Family or Intergenerational Conflicts achieved higher scores (both 4.29). This underscores the importance of contextual relevance and insightful content in enhancing perceived utility.
@@
Looking to Armando chatbot, the RAG evaluation showed clear advantages in content specificity and contextual accuracy. When responding to specific queries (e.g., ”Where is the fire happening?”), RAG-enhanced responses consistently outperformed non-RAG responses by providing accurate, contextually relevant, and actionable information. This demonstrates the crucial role of verified external knowledge integration in emergency communication scenarios.
However, the query ”Which number should I call?” highlighted challenges in clearly communicating geographical specificity. Although the system correctly identified Portugal’s emergency number (112), suggesting that users confirm their location introduced unnecessary ambiguity regarding the chatbot’s geographical awareness. Future implementations should minimise such ambiguity by presenting clear and singular region-specific recommendations, avoiding conditional or uncertain expressions.
@@ 7 Conclusion and Future Work
Further improvements should prioritise enhancing user interaction by making the visualisation of the emotional reasoning process—comprising discussion, voting, and reasoning—optional or more discreet. In addition, prompt formulations could be adapted according to the type of situation addressed, allowing for more immersive and context-sensitive responses. Finally, in the context of emergency communication, it is crucial to establish a structured syntax and annotation protocol for RAG input data, to ensure more accurate retrieval and delivery of critical information.
Funding sources
This publication is funded by FCT – Fundação para a Ciência e Tecnologia, I.P., under the project UIDB/04524/2020.
Declaration of generative AI and AI-assisted technologies in the writing process.

### v2

@@ 5.1 Response Comparison: RAG vs No RAG
The first prompt consisted solely of the user’s query: ”Where is the fire happening?”. Responses obtained with and without RAG are presented in Table 4.
@@
Analysing Table 4, the non-RAG model lacked specific contextual information regarding the ongoing incident, resulting in a generic response containing broad emergency guidance without specifying the fire’s location. Conversely, the RAG-enhanced model delivered a precise answer, clearly identifying the fire’s location as the third floor of a residential apartment at Rua de São Bento 112, Lisbon, Portugal, including coordinates (38.7169° N, 9.1399° W), evacuation status, and detailed safety instructions. This demonstrates the significant advantage of employing RAG for delivering contextually relevant and actionable information in emergency scenarios.
In the second prompt, the user asked: “A fire is happening! What should I do?” with the context “It’s difficult to breathe.”. The responses obtained are detailed in Table 5.
@@
From Table 5, both RAG and non-RAG responses effectively addressed the user’s situation by providing clear, practical evacuation and safety instructions suitable for scenarios involving breathing difficulties due to smoke inhalation.
Finally, the third prompt included the user’s query: ”Which number should I call?” with the context ”I’m alone.” and an accompanying image depicting a fire. The responses are summarised in Table 6.
@@
As observed in Table 6, the response without RAG provided generic emergency guidance and suggested calling 911 or another local emergency number without explicitly identifying the correct emergency number for Portugal. The RAG-enhanced response correctly identified the emergency number applicable in Portugal (112) but included advice for the user to verify their current country. This suggestion may inadvertently give the impression that the system is uncertain about the incident’s location, despite clearly having access to location-specific data. Therefore, while the RAG implementation successfully provided accurate and relevant information, the prompt formulation could be further refined to avoid potential confusion regarding the system’s knowledge about the user’s geographical context.
@@ 6 Discussion
In the Emotional Appropriateness category, the highest scores occurred in the themes of Job Loss or Unemployment and Difficult Personal Decisions, specifically for the question “Did the final response seem emotionally appropriate to the context of your question?”, with both themes achieving a mean score of 4.71. This indicates the system’s capability to effectively match emotional tone and content in structured or goal-focused scenarios. Conversely, responses concerning Breakups or Friendship Loss consistently yielded the lowest mean scores in empathy (4.12) and appropriateness (4.41), highlighting challenges in capturing the complexity and variability inherent to interpersonal emotional dynamics.
This trend was further supported by the alignment between user-identified emotions and those suggested by the system, evaluated via the question “Do you believe the answer(s) of the emotion(s) with the most votes were the most appropriate?”. The strongest alignment appeared in Family or Intergenerational Conflicts (mean score: 4.35), indicating effective resonance with user expectations. In contrast, Breakups or Friendship Loss scored the lowest (mean score: 3.94), suggesting a need for improved modelling of emotional subtleties in relational contexts.
Regarding Clarity and Utility, results were generally positive. Clarity received high ratings across all themes, particularly for Job Loss or Unemployment, with a maximum score of 4.71 for the question “Was the final response clear and understandable?”. However, perceived utility demonstrated greater variation. The lowest utility scores were associated with Breakups or Friendship Loss (3.88), whereas Difficult Personal Decisions and Family or Intergenerational Conflicts achieved higher scores (both 4.29). This underscores the importance of contextual relevance and insightful content in enhancing perceived utility.
@@
Looking to Armando chatbot, the RAG evaluation showed clear advantages in content specificity and contextual accuracy. When responding to specific queries (e.g., “Where is the fire happening?”), RAG-enhanced responses consistently outperformed non-RAG responses by providing accurate, contextually relevant, and actionable information. This demonstrates the crucial role of verified external knowledge integration in emergency communication scenarios.
However, the query “Which number should I call?” highlighted challenges in clearly communicating geographical specificity. Although the system correctly identified Portugal’s emergency number (112), suggesting that users confirm their location introduced unnecessary ambiguity regarding the chatbot’s geographical awareness. Future implementations should minimise such ambiguity by presenting clear and singular region-specific recommendations, avoiding conditional or uncertain expressions.
@@ 7 Conclusion and Future Work
Further improvements should prioritise enhancing user interaction by making the visualisation of the emotional reasoning process, comprising discussion, voting, and reasoning,optional or more discreet. In addition, prompt formulations could be adapted according to the type of situation addressed, allowing for more immersive and context-sensitive responses. Finally, in the context of emergency communication, it is crucial to establish a structured syntax and annotation protocol for RAG input data, to ensure more accurate retrieval and delivery of critical information.
8 Competing interests
No competing interest is declared.
9 Author contributions statement
ARO and DF conceived the study. NC and AP secured funding. ARO, GV, DF, and LF carried out the investigation. ARO, GV, DF, LF, NC, and AP developed the methodology. DF, LF, NC, and AP provided resources and supervised the work. ARO and DF developed the software. ARO, GV, DF, LF, NC, and AP validated the results. ARO, GV, DF, and LF drafted the original manuscript, and all authors revised and edited the final version.
10 Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
11 Funding
This work was supported by FCT – Fundação para a Ciência e Tecnologia, I.P., under the project UIDB/04524/2020.
12 Acknowledgments